sxdsouls:

An engineer is designing a roller coaster. The tallest peak is 310 feet high. The roller coaster travels 155 horizontal feet as it descends the hill. What is the slope of the hill? –2 –1.55 1.55 2

sxdsouls:

@snowflake0531 c':

Janiah:

what do you think the answer is

AZ:

remember slope = rise over run since we're going down from the peak, the rise will be a negative number so how high up are you? and how much do you "run"

sxdsouls:

she said 2 yeah yeah but 2 makes sense right- ._.

AZ:

no remember, we're going down so the rise is -310 and the run is 155 remember when I showed you negative slope and positive slope? doesn't the image look like it has a negative slope?

sxdsouls:

yes

AZ:

so slope = rise / run slope = -310/155 slope = ??

sxdsouls:

-2

AZ:

there you go